During the manufacture of a product, the probability that the product will be satisfactory is 90%. The products are sampled individually from a conveyor belt. Out of 5 samples, what is the probability of finding exactly 2 defective products? Use the binomial distribution table and give the exact value. Hint: A "success" in this case means a defective product.
Answer #1

Solution

given n 5, p (defective) = 1- P(satisfactory) = 1 -0.90 0.10,

g(satis factory) = 0.90

formula P(X = r) C * p *q where, C

P(exactly 2 defective products) P(X-2)

P(X 2)C 0.10 0.90-

5! P(X 2) 21(5 2*0.102 0.903

5! P(X 221 * 0.102 0.90

5 4 3! P(X=2)= 2 1) 3! 0.102 0.90

P(X 2) 100.10 0.90%

P(X =2)0.0729